# DGFusion: Depth-Guided Sensor Fusion for Robust Semantic Perception

### v1

@@ Abstract
Robust semantic perception for autonomous vehicles relies on effectively combining multiple sensors with complementary strengths and weaknesses. State-of-the-art sensor fusion approaches to semantic perception often treat sensor data uniformly across the spatial extent of the input, which hinders performance when faced with challenging conditions. By contrast, we propose a novel depth-guided multimodal fusion method that upgrades condition-aware fusion by integrating depth information. Our network, DGFusion, poses multimodal segmentation as a multi-task problem, utilizing the lidar measurements, which are typically available in outdoor sensor suites, both as one of the model’s inputs and as ground truth for learning depth. Our corresponding auxiliary depth head helps to learn depth-aware features, which are encoded into spatially varying local depth tokens that condition our attentive cross-modal fusion. Together with a global condition token, these local depth tokens dynamically adapt sensor fusion to the spatially varying reliability of each sensor across the scene, which largely depends on depth. In addition, we propose a robust loss for our depth, which is essential for learning from lidar inputs that are typically sparse and noisy in adverse conditions. Our method achieves state-of-the-art panoptic and semantic segmentation performance on the challenging MUSES and DELIVER datasets. Code and models will be available at https://github.com/timbroed/DGFusion
@@ I Introduction
Robust segmentation for automated driving systems demands reliable sensor fusion under any environmental condition. Conventional semantic perception pipelines that rely on a single sensor struggle in domains such as fog, rain, or low light, where the performance of individual sensors degrades. Prior work [1] has shown that conditioning cross-modal fusion for semantic perception on a global representation of the environmental condition of the scene can effectively guide fusion by dynamically adapting sensor contributions. However, while this approach models the global scene condition, it overlooks the spatial variability of the effect of this condition on the sensor measurements, which largely depends on the local depth.
@@
In practice, depth is a key factor affecting the variable reliability of each sensor in each part of the scene. For instance, lidars provide accurate returns for nearby objects but degrade rapidly at longer distances, e.g. in fog or snowfall [2, 3]. Radars, on the other hand, operate well at long distances even in adverse weather but suffer from strong noise due to multi-path effects, whereas cameras are reliable only in well-lit parts of the scene, which are in turn related to the 3D structure of the scene. In other words, environmental conditions introduce spatial variations in the signal-to-noise ratio of each sensor, which call for a fine-grained local adaptivity of the fusion strategy. Consequently, incorporating explicit depth cues into the fusion mechanism for semantic perception could enable such spatial adaptivity to sensor reliability, which would in turn enhance segmentation robustness in challenging environments. This hypothesis is reinforced by several works on RGB-only semantics and depth multi-tasking [4, 5], which improve segmentation performance over semantics-only learning. Despite this promise for depth-informed fusion, most state-of-the-art multi-sensor segmentation methods [6, 7, 8, 9] treat lidar solely as an input modality, overlooking its potential for improving segmentation via multi-tasking by leveraging it also as sparse ground truth for learning depth together with semantics.
@@
We fill in this gap and propose DGFusion, a depth-guided multimodal fusion network which upgrades condition-aware sensor fusion for semantic perception with explicit local depth guidance (cf. Fig. 1). In DGFusion, we introduce an auxiliary depth prediction head to enrich local features with implicit depth cues, while a global condition token is generated to guide each local-window cross-attention fusion. This dual encoding of global environmental conditions and spatially-varying depth cues enables the network to make more precise decisions based on which sensor to rely more on for each feature. By incorporating depth cues, the fusion mechanism can adaptively adjust the weighting of each modality per image region, favoring the per-case more reliable sensor. Moreover, since the noisy depth reference is already available as one of our network’s inputs, i.e. the lidar, depth supervision comes at no extra cost compared to existing fusion pipelines. The synergy between the global condition and local depth representation is especially beneficial in adverse conditions, which have been found [10] to cause pronounced spatial variations in sensor reliability.
@@
Besides, training our model on adverse conditions poses additional challenges. For example, in fog or snowfall the lidar measurements, which serve as ground truth for depth, become inherently noisy [2, 3], making the direct integration of standard depth estimation approaches and losses in our multi-tasking model infeasible. In order to get a reliable training signal across all conditions, we carefully design our depth loss to be robust to sparse and noisy ground truth by combining an outlier-robust $L_{1}$ loss with strategies commonly used in monocular depth estimation. These strategies consist of an edge-aware smoothness loss to penalize depth gradients in uniform regions of the RGB image and a novel panoptic-edge-aware smoothness loss that better captures depth discontinuities between instances of the same semantic classes compared to previous semantic-edge-aware smoothness losses.
@@
Casting the training of DGFusion as a multi-tasking problem leverages the synergy between depth and semantics with minimal changes to the inference pipeline. Extensive experiments on MUSES [11] and DeLiVER [7] demonstrate that DGFusion sets the new state of the art in multi-sensor semantic perception across diverse conditions.
@@
  • We leverage the internal dense depth features to perform a spatially-variant conditioning of our cross-modal attentive fusion on local depth tokens. Combined with a global condition token, this makes the fusion locally adaptive to the effect of this environmental condition on the sensor at the respective feature location.
  • We design an outlier-resistant depth loss suitable for sparse, noisy lidar ground truth, incorporating edge-aware smoothness and a novel panoptic-aware smoothness term for reliable training.
  • We extensively validate and ablate our network on two widely used outdoor datasets, MUSES and DeLiVER, and we set the new state of the art on multi-sensor semantic perception.
@@ II Related Work
Semantic perception encompasses tasks like semantic [12], panoptic [13], and instance [14] segmentation, which are pivotal for robotics applications such as automated driving [15]. Recent mask-based networks have advanced the state of the art [16], and unifying these tasks within a single model has proven beneficial [17]. In contrast to these RGB-only methods, our work focuses on multimodal semantic perception, leveraging sensor fusion to enhance robustness in challenging conditions where RGB data is unreliable.
@@
Sensor fusion is central in practical vision applications, as different sensors feature complementary characteristics. The availability of large multimodal datasets, from early ones like KITTI [18] to recent ones focusing on adverse conditions [7, 10, 11], has driven the development of fusion methods. Architectures have evolved from two-modality systems [19] to modular, attention-based networks that handle an arbitrary number of inputs [6, 7]. Recent approaches have integrated large-scale pre-trained models [8], explored modality-agnostic designs with shared backbones [20], and combined intra-modal with inter-modal attention [9]. The MUSES method [11] extends [21] to the multimodal setting, solving panoptic and semantic segmentation via local-window cross-attention. However, all these works fuse sensors uniformly, lacking explicit adaptation to the conditions that affect each image region non-uniformly, largely depending on depth.
@@
Condition-aware fusion utilizes information about the environmental conditions to guide perception. [19] addresses object detection by implicitly assigning larger weights to the modality that yields better detection features at late fusion. Knowledge distillation from vision-language models is used in [22] for learning modality-agnostic representations. The most closely related work to ours in this regard is [1], which guides cross-modal attention with a global camera-based condition token that is learned contrastively based on training-time supervision with image-level verbal condition descriptions. In contrast, we aim at a more fine-grained, pixel-level adaptivity to the environmental condition, as its effects on sensor data vary non-uniformly with depth.
@@
Semantics and depth are strongly correlated tasks, and it is well-established that they benefit from being learned jointly in a multi-task setup [4, 5]; cf. [23] for a comprehensive overview of such semantics-depth multi-tasking methods. Prior work has leveraged this synergy, with semantics guiding monocular depth estimation [24, 25]. Similarly, depth supervision has benefited from edge-aware smoothness losses based on image gradients [26] and semantics-aware losses that use semantic annotations [27], which we extend to use more fine-grained panoptic annotations. In contrast to previous works, we explore the use of depth in multimodal settings as a readily available input signal, e.g. from sparse lidar measurements in typical outdoor multi-sensor suites, in order to benefit semantics. While the aforementioned multi-sensor semantic perception works utilize such depth information only as input to the fusion architecture, we instead leverage it as a reference output during training, reframing sensor fusion as a multi-task problem where depth completion serves as an auxiliary task to learn informative, localized features for semantic perception. The internal depth features learned from this auxiliary task are leveraged as a spatially varying conditioning in our attentive fusion together with global condition information. This combination allows DGFusion to adaptively weigh modalities not only per-input, but also per-region, based on how depth modulates the effect of the environmental condition on each sensor.
@@ III-A Architecture
Our goal is to achieve robust multimodal segmentation in all weather conditions. We build on the state-of-the-art CAFuser [1] method, which uses global environmental condition encodings to guide the sensor fusion for the OneFormer [17] segmentation head. Following CAFuser, we project the secondary modalities (lidar, radar, and event camera) as 3-channel images onto the RGB camera plane. To further align these projected inputs with the RGB modality, we dilate the respective images with a square kernel of size $K_{\text{sensor}}$, which compensates for the sparsity of the secondary modalities and yields denser input images. In contrast to CAFuser, we treat lidar not only as an input modality but also as a source of sparse, noisy ground-truth depth. Training our network via multi-tasking semantics and depth provides robust supervision to enhance semantic features and allows us to leverage depth-informed features for locally adaptive depth-guided fusion.
@@
In Fig. 2, our complete DGFusion architecture is depicted. On the left, an arbitrary number of sensor modalities, including camera as the primary modality and otherwise depending on the dataset at hand, are first processed by shared backbones equipped with modality-specific feature adapters, resulting in per-level feature maps. These features are then split into three branches: the depth estimation branch (top), the segmentation branch (center), and the condition representation branch introduced by CAFuser [1] (bottom). The depth branch uses raw, undilated lidar ground truth to supervise depth predictions and induce depth-informed features via a lightweight fusion module, while the condition branch generates a global condition token (CT) from the RGB input. Both of the latter branches provide supplementary cues to the central depth-guided fusion module in the segmentation branch.
Depth estimation branch. At each feature level $l$, we fuse all modalities with a light-weight fusion module to produce a “depth” feature pyramid $\{d_{l}\}_{l=1}^{4}$. In our fusion module, we concatenate all modality features along the channel dimension and process the result with a 2-layer multilayer perceptron (MLP) with a bottleneck, followed by a residual connection from the RGB features to ensure stable gradient flow. The resulting depth features are passed to our auxiliary depth head and supervised by our depth loss described in Sec. III-B. Note that we fuse all input modalities to exploit all available information and to fully utilize their complementary strengths. We use a Semantic FPN [13] decoder as our auxiliary depth head. This head upsamples the features $d_{l}$ from each level of the pyramid with multiple convolutions and bilinear upsampling steps until it reaches 1/4 scale. The respective outputs from the various levels are then summed and finally transformed into a single depth prediction. This structure fuses coarse and fine information efficiently, enabling robust depth estimation at various scales. At inference time, we drop the auxiliary depth head, but keep the depth features $d_{l}$ and still feed them to our fusion module. This setup keeps the alterations to the inference pipeline to a minimum while still computing depth-informed features for guiding fusion in the semantic branch in a locally adaptive fashion.
@@
Condition representation branch. We adopt the CAFuser [1] strategy to produce a global CT. Specifically, we flatten the highest-level features from the RGB camera and feed them into a Transformer composed of 2 encoder and 2 decoder layers. The resulting token is then supervised during training with a verbo-visual contrastive loss [1, 17, 28] that leverages text prompts describing the environmental conditions in detail.
@@
Depth-guided fusion. We use the depth features $\{d_{l}\}_{l=1}^{4}$ to guide sensor fusion in our main semantic branch. Our Depth-Guided Fusion module follows CAFuser in utilizing parallel multi-window cross-attention fusion [6] to fuse in parallel each secondary modality via cross-attention with the primary camera modality within $K_{w}\times K_{w}$ local windows. This fusion design is efficient and scales well with the number of input modalities. In each Depth-Guided Fusion module, the RGB features, the features from the relevant secondary modality and the depth features are split into local windows of size $K_{w}\times K_{w}$. The depth features from each local window are passed through a 2D convolution and a mean pooling to yield a local Depth Token (DT), which represents the local depth context per window. This local DT is concatenated with the corresponding RGB tokens and the global CT to form the set of queries for the cross-attention operation with the secondary modality. After cross-attention, the DT and CT are removed to restore the original spatial dimensions of the RGB tokens. Finally, the fused features from all parallel Depth-Guided Fusion modules are summed with the residual features and passed to the segmentation head, which outputs the panoptic prediction. This design ensures that the predicted depth assists the multi-sensor semantic fusion while imposing minimal changes to the overall architecture. Crucially, the depth head is only active at training time, allowing the network to learn richer feature representations via multi-task supervision. At inference, we do not require the depth head, as the Depth-Guided Fusion module only needs the depth features that are computed from the sensor inputs. Each local-window cross-attentive fusion is jointly guided both by the global condition context encoded in the CT and the localized depth cues encoded in the respective local DT, which together model well the local effect of the present condition on the reliability of each sensor and adapt fusion properly.

### v2

@@ Abstract
Robust semantic perception for autonomous vehicles relies on effectively combining multiple sensors with complementary strengths and weaknesses. State-of-the-art sensor fusion approaches to semantic perception often treat sensor data uniformly across the spatial extent of the input, which hinders performance when faced with challenging conditions. By contrast, we propose a novel depth-guided multimodal fusion method that upgrades condition-aware fusion by integrating depth information. Our network, DGFusion, poses multimodal segmentation as a multi-task problem, utilizing the lidar measurements, which are typically available in outdoor sensor suites, both as one of the model’s inputs and as ground truth for learning depth. Our corresponding auxiliary depth head helps to learn depth-aware features, which are encoded into spatially varying local depth tokens that condition our attentive cross-modal fusion. Together with a global condition token, these local depth tokens dynamically adapt sensor fusion to the spatially varying reliability of each sensor across the scene, which largely depends on depth. In addition, we propose a robust loss for our depth, which is essential for learning from lidar inputs that are typically sparse and noisy in adverse conditions. Our method achieves state-of-the-art panoptic and semantic segmentation performance on the challenging MUSES and DeLiVER datasets. Code and models are available at https://github.com/timbroed/DGFusion
@@ I Introduction
Robust segmentation for automated driving systems demands reliable sensor fusion under any environmental condition. Conventional semantic perception pipelines that rely on a single sensor struggle in domains such as fog, rain, or low light, where the performance of individual sensors degrades. Prior work [cafuser] has shown that conditioning cross-modal fusion on a global Condition Token (CT) that encodes the scene-wide environmental condition can effectively guide fusion by dynamically adapting sensor contributions. However, while this global approach models the overall scene condition, it overlooks the spatial variability of the effect of this condition on the sensor measurements, which largely depends on the local depth.
@@
In practice, depth is a key factor affecting the variable reliability of each sensor in each part of the scene. For instance, lidars provide accurate returns for nearby objects but degrade rapidly at longer distances, e.g. in fog or snowfall [hahner2021fog, hahner2022lidar]. Radars, on the other hand, operate well at long distances even in adverse weather but suffer from strong noise due to multi-path effects, whereas cameras are reliable only in well-lit parts of the scene, which are in turn related to the 3D structure of the scene. In other words, environmental conditions introduce spatial variations in the signal-to-noise ratio of each sensor, which call for a fine-grained local adaptivity of the fusion strategy. Consequently, incorporating explicit depth cues into the fusion mechanism for semantic perception could enable such spatial adaptivity to sensor reliability, which would in turn enhance segmentation robustness in challenging environments. This hypothesis is reinforced by several works on RGB-only semantics and depth multi-tasking [bruggemann2021exploring, rahman2024semi], which improve segmentation performance over semantics-only learning. Despite this promise for depth-informed fusion, most state-of-the-art multi-sensor segmentation methods [broedermann2023hrfuser, zhang2023delivering, li2024stitchfusion, jia2024geminifusion] treat lidar solely as an input modality, overlooking its potential for improving segmentation via multi-tasking by leveraging its pixel-aligned sparse depth measurements as ground truth for learning depth together with semantics.
@@
We fill in this gap and propose DGFusion, a depth-guided multimodal fusion network which upgrades condition-aware sensor fusion for semantic perception with explicit local depth guidance (cf. Fig. 1). Building on CAFuser [cafuser], we adopt its shared backbone with modality-specific feature adapters and its global CT that encodes scene-wide environmental conditions. Beyond this, in DGFusion we introduce an auxiliary depth prediction head that leverages lidar’s pixel-aligned sparse depth measurements as noisy supervision to enrich local features with implicit depth cues to guide each local-window cross-attention fusion. This dual encoding of global environmental conditions and spatially-varying depth cues enables the network to make more precise decisions based on which sensor to rely more on for each feature. By incorporating depth cues, the fusion mechanism can adaptively adjust the weighting of each modality per image region, favoring the per-case more reliable sensor. Moreover, since these noisy depth measurements are already available as one of our network’s inputs, i.e. the lidar, depth supervision comes at no extra cost compared to existing fusion pipelines. The synergy between the global condition and local depth representation is especially beneficial in adverse conditions, which have been found [seeing_through_fog] to cause pronounced spatial variations in sensor reliability.
@@
Besides, training our model on adverse conditions poses additional challenges. For example, in fog or snowfall the lidar measurements, which provide pixel-aligned depth supervision, become inherently noisy [hahner2021fog, hahner2022lidar], making the direct integration of standard depth estimation approaches and losses in our multi-tasking model infeasible. In order to get a reliable training signal across all conditions, we carefully design our depth loss to be robust to sparse and noisy ground truth by combining an outlier-robust $L_{1}$ loss with strategies commonly used in monocular depth estimation. These strategies consist of an edge-aware smoothness loss to penalize depth gradients in uniform regions of the RGB image and a novel panoptic-edge-aware smoothness loss that better captures depth discontinuities between instances of the same semantic classes compared to previous semantic-edge-aware smoothness losses.
@@
Casting the training of DGFusion as a multi-tasking problem leverages the synergy between depth and semantics with minimal changes to the inference pipeline. Extensive experiments on MUSES [MUSES] and DeLiVER [zhang2023delivering] demonstrate that DGFusion sets the new state of the art in multi-sensor semantic perception across diverse conditions.
@@
  • We leverage the internal depth features to perform a spatially-variant conditioning of our cross-modal attentive fusion on local depth tokens. Combined with a global CT, this makes our fusion locally adaptive to the effect of this condition on the sensor at the respective feature location.
  • We design an outlier-resistant depth loss on sparse, noisy lidar ground truth, incorporating edge-aware and panoptic-aware smoothness priors for reliable training.
  • We extensively validate and ablate DGFusion on two widely used outdoor datasets, MUSES and DeLiVER, and set the new state of the art on multi-sensor segmentation.
@@ II Related Work
Semantic perception encompasses tasks like semantic [Cityscapes], panoptic [kirillov2019panoptic], and instance [he2017mask] segmentation, which are pivotal for robotics applications such as automated driving [does:vision:matter:for:action]. Recent mask-based networks have advanced the state of the art [li2023mask], and unifying these tasks within a single model has proven beneficial [jain2023oneformer]. In contrast to these RGB-only methods, our work focuses on multimodal semantic perception, leveraging sensor fusion to enhance robustness in challenging conditions where RGB data is unreliable.
@@
Sensor fusion is central in practical vision applications, as different sensors feature complementary characteristics. The availability of large multimodal datasets, from early ones like KITTI [kitti] to recent ones focusing on adverse conditions [zhang2023delivering, seeing_through_fog, MUSES], has driven the development of fusion methods. Architectures have evolved from two-modality systems [chaturvedi_pay] to modular, attention-based networks that handle an arbitrary number of inputs [broedermann2023hrfuser, zhang2023delivering]. Recent approaches have integrated large-scale pre-trained models [li2024stitchfusion], explored modality-agnostic designs with shared backbones [centeringthevalue], and combined intra-modal with inter-modal attention [jia2024geminifusion]. The MUSES method [MUSES] extends [cheng2022masked] to the multimodal setting, solving panoptic and semantic segmentation via local-window cross-attention. However, all these works fuse sensors uniformly, lacking explicit adaptation to the conditions that affect each image region non-uniformly, largely depending on depth.
@@
Condition-aware fusion utilizes information about the environmental conditions to guide perception. [chaturvedi_pay] addresses object detection by implicitly assigning larger weights to the modality that yields better detection features at late fusion. Knowledge distillation from vision-language models is used in [learningmodalityagnostic] for learning modality-agnostic representations. The most closely related work to ours in this regard is [cafuser], which guides cross-modal attention with a global camera-based CT that is learned contrastively based on training-time supervision with image-level verbal condition descriptions. In contrast, we aim at a more fine-grained, pixel-level adaptivity to the environmental condition, as its effects on sensor data vary non-uniformly with depth.
@@
Semantics and depth are strongly correlated tasks, and it is well-established that they benefit from being learned jointly in a multi-task setup [bruggemann2021exploring, rahman2024semi]; cf. [vandenhende2021multitask] for a comprehensive overview of such semantics-depth multi-tasking methods. Prior work has leveraged this synergy, with semantics guiding monocular depth estimation [chen2019towards, zhu2020edge]. Similarly, depth supervision has benefited from edge-aware smoothness losses based on image gradients [godard2017unsupervised] and semantics-aware losses that use semantic annotations [li2024semantic], which we extend to use more fine-grained panoptic annotations. In contrast to previous works, we explore the use of depth in multimodal settings as a readily available input signal, e.g. from sparse lidar measurements in typical outdoor multi-sensor suites, in order to benefit semantics. While the aforementioned multi-sensor semantic perception works utilize such depth information only as input to the fusion architecture, we instead leverage it as a reference output during training, reframing sensor fusion as a multi-task problem where depth completion serves as an auxiliary task to learn informative, localized features for semantic perception. The internal depth features learned from this auxiliary task are leveraged as a spatially varying conditioning in our attentive fusion together with global condition information. This combination allows DGFusion to adaptively weigh modalities not only per-input, but also per-region, based on how depth modulates the effect of the environmental condition on each sensor.
@@ III-A Architecture
Our goal is to achieve robust multimodal segmentation in all weather conditions. We build on CAFuser [cafuser], which processes all modalities through a shared backbone with lightweight modality-specific feature adapters to align heterogeneous inputs into a common feature space, and uses global environmental condition encodings to guide the sensor fusion for the OneFormer [jain2023oneformer] segmentation head. Following CAFuser, we project the secondary modalities (lidar, radar, and event camera) as 3-channel images onto the RGB camera plane and dilate them with a kernel of size $K_{\text{sensor}}$ to compensate for sparsity. We adopt CAFuser’s shared backbone with feature adapters and its global condition token (CT). In contrast to CAFuser, we treat lidar not only as an input modality but also as sparse, noisy ground-truth depth for multi-task learning, which poses additional challenges in adverse conditions where lidar measurements degrade. With our carefully designed robust losses (cf. Sec. III-B), this enables our network to learn depth-informed features that provide additional supervision for semantics and enable locally adaptive depth-guided fusion via local depth tokens.
@@
In Fig. 2, our complete DGFusion architecture is depicted. On the left, an arbitrary number of sensor modalities are processed by the shared backbone with modality-specific feature adapters, resulting in per-level feature maps. These features are then split into three branches: the depth estimation branch (top), the segmentation branch (center), and the condition representation branch adopted from CAFuser (bottom). The depth branch uses raw, undilated lidar ground truth to supervise depth predictions and induce depth-informed features via a lightweight fusion module, while the condition branch generates a global CT from the RGB input. Both branches provide supplementary cues to the central depth-guided fusion module.
Depth estimation branch. At each feature level $l\in\{1,2,3,4\}$, we fuse all modalities with a light-weight fusion module to produce a “depth” feature pyramid $\mathbf{d}_{l}$. Given RGB features $\mathbf{F}_{\text{rgb}}^{l}$ and secondary modality features $\mathbf{F}_{\text{sec}}^{l,m}$ for modality $m\in\{1,\ldots,M\}$, we compute
$\mathbf{d}_{l}=\text{MLP}([\mathbf{F}_{\text{rgb}}^{l},\mathbf{F}_{\text{sec}}^{l,1},\ldots,\mathbf{F}_{\text{sec}}^{l,M}])+\mathbf{F}_{\text{rgb}}^{l},$ (1)
where $\cdot$] denotes concatenation and MLP is a 2-layer perceptron with a bottleneck. The residual connection from RGB features ensures stable gradient flow. The resulting depth features are passed to our auxiliary depth head and supervised by our depth loss described in Sec. III-B. Note that we use all input modalities to exploit all available information and to fully utilize their complementary strengths.
We use a Semantic FPN [kirillov2019panoptic] decoder as our auxiliary depth head, which upsamples each depth feature $\mathbf{d}_{l}$ to 1/4 resolution and aggregates them:
$\hat{D}=\sum_{l=1}^{4}\text{Upsample}(\mathbf{d}_{l}),$ (2)
where Upsample consists of multiple convolutions and bilinear upsampling steps. This structure fuses coarse and fine information efficiently, enabling robust depth estimation at various scales. At inference, we drop the auxiliary depth head but keep the depth features $\mathbf{d}_{l}$ for guiding fusion in the semantic branch.
@@
Condition representation branch. We adopt CAFuser’s [cafuser] strategy to produce a global CT. We flatten the highest-level RGB features and process them with a lightweight Transformer composed of 2 encoder and 2 decoder layers:
$\mathbf{t}_{c}=\text{Transformer}(\text{Flatten}(\mathbf{F}_{\text{rgb}}^{4})),$ (3)
supervised during training with a verbo-visual contrastive loss [cafuser, jain2023oneformer, bruggemann2023contrastive] that leverages text prompts describing the environmental conditions in detail.
Depth-guided fusion. We use the depth features $\mathbf{d}_{l}$ to guide sensor fusion in our main semantic branch. Our Depth-Guided Fusion module utilizes parallel multi-window cross-attention fusion [broedermann2023hrfuser] to fuse in parallel each secondary modality via cross-attention with the primary camera modality within $K_{w}\times K_{w}$ local windows. This fusion design is efficient and scales well with the number of input modalities. For each level $l$ and secondary modality $m$, we partition RGB features $\mathbf{F}_{\text{rgb}}^{l}$, secondary features $\mathbf{F}_{\text{sec}}^{l,m}$, and depth features $\mathbf{d}_{l}$ into $I_{l}$ windows of size $K_{w}\times K_{w}$. For each window $i\in\{1,\ldots,I_{l}\}$, we extract a local Depth Token (DT):
$\mathbf{t}_{d}^{l,i}=\text{Pool}_{\text{mean}}(\text{Conv}(\mathbf{d}_{l}^{i})),$ (4)
which encodes local depth context. We map the CT to the feature dimensionality of level $l$ with a fully connected layer:
$\mathbf{t}_{c}^{l}=\text{FC}(\mathbf{t}_{c}).$ (5)
We then form query features by concatenating RGB tokens with the global CT $\mathbf{t}_{c}^{l}$ and the local DTs $\mathbf{t}_{d}^{l,i}$ into $\mathbf{F}_{q}^{l,i}=[\mathbf{F}_{\text{rgb}}^{l,i},\mathbf{t}_{c}^{l},\mathbf{t}_{d}^{l,i}]$. Depth-guided fusion is performed locally by letting the condition- and depth-aware RGB features cross-attend to each secondary modality $m$ in window $i$:
$\mathbf{F}_{\text{fused}}^{l,m,i}=\mathbf{F}_{q}^{l,i}+\text{Softmax}(\mathbf{Q}_{l,i}\mathbf{K}_{l,m,i}^{T})\mathbf{V}_{l,m,i},$ (6)
where the queries are computed as linear projections of the enriched RGB features $\mathbf{Q}_{l,i}=\mathbf{W}_{q}^{l,i}\mathbf{F}_{q}^{l,i}$, and respectively for the keys and values and the secondary modality features via $\mathbf{K}_{l,m,i}=\mathbf{W}_{k}^{l,m,i}\mathbf{F}_{\text{sec}}^{l,m,i}$ and $\mathbf{V}_{l,m,i}=\mathbf{W}_{v}^{l,m,i}\mathbf{F}_{\text{sec}}^{l,m,i}$. After the cross-attention layer of (6), we remove the last two tokens from $\mathbf{F}_{\text{fused}}^{l,m,i}$, which correspond to the CT and the local DT, thus aligning the spatial dimensions of the resulting features $\hat{\mathbf{F}}_{\text{fused}}^{l,m,i}$ back to those of $\mathbf{F}_{\text{rgb}}^{l,i}$. We reassemble windows and sum over all modalities with a residual connection:
$\mathbf{F}_{\text{out}}^{l}=\sum_{m=1}^{M}\text{Reassemble}\left(\{\hat{\mathbf{F}}_{\text{fused}}^{l,m,i}\}_{i=1}^{I_{l}}\right)+\mathbf{F}_{rgb}^{l}.$ (7)
The resulting feature pyramid $\{\mathbf{F}_{\text{out}}^{l}\}_{l=1}^{4}$ is passed to the segmentation head, which outputs the panoptic prediction.
@@
This design ensures that the predicted depth assists the multi-sensor semantic fusion while imposing minimal changes to the overall architecture. Crucially, the depth head is only active at training time, allowing the network to learn richer feature representations via multi-task supervision. At inference, we do not require the depth head, as the Depth-Guided Fusion module only needs the depth features that are computed from the sensor inputs. Each local-window cross-attentive fusion is jointly guided both by the global condition context encoded in the CT and the localized depth cues encoded in the respective local DT, which together model well the local effect of the present condition on the reliability of each sensor and adapt fusion properly.